# imitation learning BEV bird’s-eye view

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
After years of progress, autonomous driving systems are able to handle increasingly complex situations [1]. This is enabled, in part, by solving several aspects of driving with learned modules, such as perception [2, 3, 4] and motion forecasting [5, 6]. Recently, there has been increased interest in managing the complexity of human behavior in traffic by also learning the planning task [7, 8, 9], which has been accelerated through the emergence of motion forecasting- and planning-centric benchmarks and datasets [10, 11]. Most of this work assumes a simplified setting where the ground-truth world state is available as input and focuses on accuracy of the planned trajectory with respect to human driving. As a result, robustness to the residual risk of errors in the perception system, which is ultimately just an imperfect model operating on incomplete observations of the world, remains underexplored.
@@
In this work, we aim to illuminate the susceptibility of learned planning, which is often brittle in the face of o.o.d inputs [12], to these errors. Recent seminal studies [13, 14] have approached this problem from the lens of adversarial attacks. To this end, these methods first construct a perception error model (PEM[15, 16], which allows sampling multiple noisy estimates imitating a target 3D object detector given a ground-truth scene representation as context. Then, by leveraging the PEM as a proxy of the detector, challenging samples that stress-test the target planner can be found by employing an adversarial search strategy. While promising, these works consider simple, synthetic scenes that do not capture the complexities of real-world data. Moreover, they employ simple PEMs that phrase the error modeling task as an isolated, per-object perturbation of the ground-truth state, rather than jointly reasoning over the entire scene context. Hence, they cannot faithfully model the intricacies of the error patterns exhibited by modern 3D object detectors, such as duplicate detections resulting in false-positives and correlations in errors for groups of objects.
Motivated by this, we propose Emperror, a novel generative empirical error model based on the transformer architecture, that can more faithfully capture the error characteristics of a target detector. Our key idea is to leverage the attention mechanism and a flexible set of latent queries to model the full range of failure modes, including false-positives, in a scene-consistent manner. Furthermore, Emperror provides a prior over different error patterns for a given scene context, enabling us to draw adversarial, yet plausible samples to probe the robustness of a target planner. Building on these advantages, we design a framework to probe the robustness of learned planners to noisy perception inputs, which is visualized in Fig. 1. We then apply the proposed framework to an imitation learning (IL)-based planner, modeling three different modern camera-based 3D object detectors, and show learned planning is indeed vulnerable to plausible noise from the long-tail of perception errors. We believe Emperror can serve as a valuable tool for data-driven evaluation of self-driving planners.
Contributions: (1) We propose Emperror, a novel transformer-based, generative PEM for probing planning, that can more faithfully imitate modern object detectors than previous work. (2) We integrate Emperror into a framework for probing the robustness of a planner to noise in its perception system. (3) We show that our proposed PEM can imitate modern 3D object detectors and enables the generation of challenging yet plausible errors that can be used to evaluate planning in a data-driven manner.
@@ II Related Work
Perception Error Models: Accurately modeling the noise characteristics of a perception module enables an analysis of typical failure modes and informs the design of downstream modules. It also enables efficient simulation for modular autonomy systems without the challenges of a visual domain gap [17, 16] by using a perception error model (PEM) as a proxy of the perception module. Several works [18, 19] rely on classical statistical models often coupled with elements of manual design to capture regression errors and false-negatives in detection tasks. While these approaches yield lightweight, easily interpretable PEMs, they are limited in the fidelity and complexity of noise patterns that can be modeled. Recent work addresses this limitation by instantiating the PEM through a neural network. In [13], the authors use a feed-forward network to model false-negative detections. In [16, 14], the authors model spatial errors and false-negatives using a probabilistic feed-forward network to enable efficient evaluation of decision making modules in simulation. In [17, 20], the authors propose fully-convolutional PEMs that mimick the error characteristics of LiDAR-based detectors [21, 22]. As they resemble the overall architecture of a standard convolutional one-stage detector, they can naturally model regression errors, false-negatives and false positives but are not probabilistic and thus do not permit sampling. In summary, the PEMs proposed in previous work either do not fully capture the error characteristics of a target perception system, do not permit sampling, or both. As these are prerequisites to find plausible worst-case perception errors for planning, we propose a novel PEM that satisfies these requirements.
Generating Safety-Critical Scenarios: When deployed, self-driving systems are required to robustly handle rare and potentially safety-critical scenarios from the long-tail of driving. Since real-world data collection is limited in scalability and diversity, there has been increased interest in the automated generation of safety-critical scenarios in recent years. The majority of this work focuses on automatically altering the behavior of other traffic participants to induce failure in the target autonomy system [23, 24, 25, 26]. Rather than testing against external long-tail behavior, we instead look inward to examine the effects of long-tail noise in the autonomy system’s own perception system by sampling from a PEM. While this high-level idea has been explored before, previous work [13, 14] uses simple PEMs that do not fully capture all dependencies and failure modes, and tests in simple, synthetic scenes that do not reflect the complexity of real-world data. In contrast, we propose a novel transformer-based PEM that fully models the false-positive, false-negative and regression error characteristics of a given 3D-object detector and apply our framework on challenging real-world data.
@@ III Method
Problem Statement: We are interested in stress-testing modular autonomy systems, which we will consider to consist of a 3D object detector as the perception module and a downstream planner for purposes of this study. Let us denote the true world state as a set of vectors $\mathbf{S}=\{\mathbf{s}_{0},\cdots,\mathbf{s}_{N}\}\in\mathbb{R}^{d_{s}}$ describing the position, heading angle, spatial dimensions, first- and second-order longitudinal- and angular dynamics in the ego vehicle’s frame as well as the semantic class for each of $N$ objects in the scene. The object detector then processes raw sensor observations of the world into a set of 3D bounding boxes $\hat{\mathbf{B}}=\{\hat{\mathbf{b}}_{0},\cdots,\hat{\mathbf{b}}_{\hat{N}}\}\in\mathbb{R}^{d_{b}}$ similar to $\mathbf{S}$, but with the dynamics reduced to a velocity vector. Based on the 3D bounding boxes, as well as a rasterized bird’s-eye view (BEV) map $\mathcal{M}\in\mathcal{R}^{h\times w\times 5}$ containing information on the static scene layout, the planner then computes a future trajectory $\boldsymbol{\tau}$ to be driven.
Due to the incompleteness of the observations and approximate nature of the model, the detector incurs errors in the form of false-positives, false-negatives and inaccurate regression of bounding box parameters when compared to the ground-truth set derived directly from $\mathbf{S}$. To probe the planner’s sensitivity to imperfect perception, we would like to find worst-case, but plausible errors for a given scene that drive the planned trajectory towards violation of safety constraints. To this end, we assume the perception errors follow a certain distribution which is characteristic for the specific detection model, and learn a conditional generative PEM as a proxy that imitates the detector given the ground-truth scene context $(\mathbf{S},\mathcal{M})$ as input. By repeatedly drawing samples $\tilde{\mathbf{B}}=\{\tilde{\mathbf{b}}_{0},\cdots,\tilde{\mathbf{b}}_{\tilde{N}}\}\in\mathbb{R}^{d_{b}}$ from this proxy and measuring the quality of the corresponding planned trajectory, we aim to gauge the influence of noise patterns and optimize for planning failures, such as a collision.
@@ III-A Conditional Generative PEM
We cast the imitation of perception errors given a ground-truth scene state as a set-to-set modelling problem. For this, we propose a conditional generative model building on a transformer encoder-decoder architecture [27] under the framework of CVAEs [28]. Specifically, we aim to construct a latent variable model capturing
@@
where $\mathbf{Z}=\{\mathbf{z}_{0},\cdots,\mathbf{z}_{\hat{N}}\}\in\mathbb{R}^{d_{z}}$ is a set of per-object latent variables of dimensionality $d_{z}$ explaining the stochasticity of the target detector’s perception errors by capturing different noise patterns. This allows sampling multiple plausible sets of detections $\tilde{\mathbf{B}}\sim p_{\theta}(\hat{\mathbf{B}}\mid\mathbf{S},\mathcal{M})$ by applying different latent error characteristics sampled from a learned conditional Gaussian prior modeling $\mathbf{Z}\sim p_{\theta}(\mathbf{Z}\mid\mathbf{S},\mathcal{M})$ to a given scene context $(\mathbf{S},\mathcal{M})$. This is done via a deterministic decoder $\hat{\mathbf{B}}=f_{\theta}(\mathbf{Z},\mathbf{S},\mathcal{M})$. As evaluating the integral in (1) is intractable, learning under the CVAE framework utilizes a learned approximation to the posterior $q_{\phi}(\mathbf{Z}\mid\mathbf{S},\mathcal{M},\hat{\mathbf{B}})$ to maximize the evidence lower bound (ELBO[28]:
@@
Although both the prior and the posterior distribution are factorized over objects, the distribution over each latent variable $\mathbf{z}_{n}\in\mathbf{Z}$ is computed from the full set of ground-truth states $\mathbf{S}$ (and detector outputs $\hat{\mathbf{B}}$ for the posterior). Similarly, the decoder considers the joint sets of per-object latent variables $\mathbf{Z}$ and states $\mathbf{S}$ to output individual detections $\tilde{\mathbf{b}}_{n}\in\tilde{\mathbf{B}}$. This enables our model to generate realistic, scene-consistent noise patterns by allowing it to capture higher-order relationships. In the following we briefly describe the main components in our architecture, as depicted in Fig. 2.
@@
Decoder: Given the set of initial hypotheses and map context, the decoder applies a sampled latent error pattern $\mathbf{Z}$ to produce a set of proxy bounding boxes $\tilde{\mathbf{B}}=f_{\theta}(\mathbf{Z},\mathbf{S},\mathcal{M})$. To this end, the information in both $\mathbf{Q}_{H}$ and $\mathbf{Z}$ is first fused by applying an MLP to the feature-wise concatenation of each pair of vectors $(\mathbf{q}_{n},\mathbf{z}_{n})$, forming $\tilde{\mathbf{Q}}_{H}\in\mathcal{R}^{d_{h}}$. Similarly to the prior encoder, $\tilde{\mathbf{Q}}_{H}$ is then refined through of a series of $L_{d}$ transformer layers performing self-attention and cross-attending to $\mathbf{Q}_{\mathcal{M}}$. Finally, the refined latent queries are decoded to the bounding box regression parameters and independent per-class sigmoid classification scores. This mirrors the prevalent output configuration used in state-of-the-art 3D detectors [2, 29, 3].
Training: We optimize the model on a dataset of tuples $(\mathbf{S},\mathcal{M},\hat{\mathbf{B}})$ collected by running inference with a target detector on the corresponding sensor data. Similar to prior work [30, 26] , we use a modified CVAE objective $\mathcal{L}_{\text{cvae}}=\mathcal{L}_{\text{recon}}+\beta\mathcal{L}_{\text{JS}}$ consisting of a reconstruction error and a divergence regularizer:
@@ III-B Imitation Learning-based Planner
We consider a simple transformer-based planner that is prototypical of the planning modules proposed in recent literature [7, 34, 35]. It operates on BEV projections of $\hat{\mathbf{B}}$ or $\tilde{\mathbf{B}}$, which are encoded via an MLP to form $\mathbf{Q}_{H}^{\pi}\in\mathbb{R}^{d_{\pi}}$, as well as a set of map features $\mathbf{Q}_{\mathcal{M}}^{\pi}\in\mathbb{R}^{d_{\pi}}$ obtained from a convolutional encoder. The planner then forms a latent plan by refining an initial constant learnable embedding $\mathbf{q}_{\text{ego}}^{\pi}$ through a series of $L_{\pi}$ transformer layers consisting of two cross-attention blocks attending to the encoded detection results $\mathbf{Q}_{H}^{\pi}$ and map features $\mathbf{Q}_{\mathcal{M}}^{\pi}$. To provide additional context, we add embeddings of the current speed of the ego vehicle $\vartheta^{\text{ego}}$, as well as a high-level navigation command $c^{\text{nav}}$ (i.e. go-straight, turn-left or turn-right) to the initial $\mathbf{q}_{\text{ego}}^{\pi}$. Finally, an MLP decodes the latent plan feature to a trajectory $\hat{\boldsymbol{\tau}}=\pi_{\omega}(\mathbf{B},\mathcal{M},\vartheta^{\text{ego}},c^{\text{nav}})\in\mathbb{R}^{T_{\text{plan}}\times 3}$ of future waypoints consisting of a BEV position and heading angle. The model is trained via standard imitation-learning on expert trajectories using an $l_{1}$-loss.

### v2

@@ Abstract
Index Terms:
Deep Learning Methods; Object Detection, Segmentation and Categorization; Autonomous Agents
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
After years of progress, autonomous driving systems are able to handle increasingly complex situations [1]. This is enabled, in part, by solving several aspects of driving with learned modules, such as perception [2, 3, 4] and motion forecasting [5, 6]. Recently, there has been increased interest in managing the complexity of human behavior in traffic by also learning the planning task [7, 8], which has been accelerated through the emergence of motion forecasting- and planning-centric benchmarks and datasets [9, 10]. Most of this work assumes a simplified setting where the ground-truth world state is available as input and focuses on accuracy of the planned trajectory with respect to human driving demonstrations. As a result, robustness to the residual risk of errors in the perception system, which is ultimately just an imperfect model operating on incomplete observations of the world, remains underexplored.
@@
In this work, we aim to illuminate the susceptibility of learned planning, which is often brittle in the face of o.o.d inputs [11], to these errors. Recent seminal studies [12, 13] have approached this problem from the lens of adversarial attacks. To this end, they first construct a perception error model (PEM[14, 15], which allows sampling multiple noisy estimates imitating a target 3D object detector given a ground-truth scene representation as context. Then, by leveraging the PEM as a proxy of the detector, challenging samples that stress-test the target planner can be found by employing an adversarial search strategy. While promising, these works consider simple, synthetic scenes that do not capture the complexities of real-world data. Moreover, they employ simple PEMs that phrase the error modeling task as an isolated, per-object perturbation of the ground-truth state, rather than jointly reasoning over the entire scene context. Hence, they cannot faithfully model the intricacies of the error patterns exhibited by modern 3D object detectors, such as duplicate detections resulting in false-positives and correlations in errors for groups of objects.
Motivated by this, we propose Emperror, a novel generative empirical error model based on the transformer architecture, that can more faithfully capture the error characteristics of a target detector. Our key idea is to leverage the attention mechanism and a flexible set of latent queries to model the full range of failure modes, including false-positives, in a scene-consistent manner. Furthermore, Emperror provides a prior over different error patterns for a given scene context, enabling us to draw adversarial, yet plausible samples to stress-test a target planner. Building on these advantages, we design a framework to probe the robustness of learned planners to noisy perception inputs, which is visualized in Fig. 1. We then apply the proposed framework to an imitation learning (IL)-based planner, modeling three different modern camera-based 3D object detectors, and show learned planning is indeed vulnerable to plausible noise from the long-tail of perception errors. We believe Emperror can serve as a valuable tool for data-driven evaluation of self-driving planners.
Contributions: (1) We propose Emperror, a novel flexible, transformer-based, generative PEM for probing planning, that can more faithfully imitate modern object detectors than previous work. (2) We propose and integrate Emperror into a framework for probing the robustness of a planner to noise in its perception system. (3) We demonstrate that this framework can reveal unsafe behavior in learned planning, even for minor, plausible detection errors.
@@ II Related Work
Perception Error Models: Accurately modeling the noise characteristics of a perception module enables an analysis of typical failure modes and informs the design of downstream modules. Several works [16, 14, 17] rely on classical statistical models often coupled with elements of manual design to capture regression errors and false-negatives in detection tasks. While these approaches yield lightweight, easily interpretable PEMs, they are limited in the fidelity and complexity of noise patterns that can be modeled. Recent work addresses this limitation by instantiating the PEM through a neural network. In [12], the authors use a feed-forward network to model false-negative detections. In [15, 13], the authors model spatial errors and false-negatives using a probabilistic feed-forward network to enable efficient evaluation of decision making modules in simulation. In [18, 19], the authors propose fully-convolutional PEMs that mimic the error characteristics of LiDAR-based detectors [20, 21]. As they resemble the overall architecture of a standard convolutional one-stage detector, they can model regression errors, false-negatives and false-positives but are not probabilistic and thus do not permit sampling. In summary, the PEMs proposed in previous work either do not fully capture the error characteristics of a target perception system, do not permit sampling, or both. As these are prerequisites to find plausible worst-case perception errors for planning, we propose a novel PEM that satisfies these requirements.
Generating Safety-Critical Scenarios: When deployed, self-driving systems are required to robustly handle rare and potentially safety-critical scenarios from the long-tail of driving. Since real-world data collection is limited in scalability and diversity, there has been increased interest in the automated generation of safety-critical scenarios in recent years. The majority of this work focuses on automatically altering the behavior of other traffic participants to induce failure in the target autonomy system [22, 23, 24, 25, 26]. Rather than testing against external long-tail behavior, we instead look inward to examine the effects of long-tail noise in the autonomy system’s own perception system by sampling from a PEM. While this high-level idea has been explored before, previous work [12, 13] uses simple PEMs that do not fully capture all dependencies and failure modes, and tests in simple, synthetic scenes that do not reflect the complexity of real-world data. In contrast, we propose a novel transformer-based PEM that fully models the false-positive, false-negative and regression error characteristics of a given 3D-object detector and apply our framework on challenging real-world data.
@@ III Method
Problem Statement: We are interested in supplementing offline evaluation by stress-testing modular autonomy systems, which we will consider to consist of a 3D object detector as the perception module and a downstream planner in this study. Let us denote the true world state as a set of vectors $\mathbf{S}=\{\mathbf{s}_{0},\cdots,\mathbf{s}_{N}\}\in\mathbb{R}^{d_{s}}$ describing the position, heading angle, spatial dimensions, first- and second-order longitudinal- and angular dynamics in the ego vehicle’s frame as well as the semantic class for each of $N$ objects in the scene. The object detector then processes raw sensor observations of the world into a set of 3D bounding boxes $\hat{\mathbf{B}}=\{\hat{\mathbf{b}}_{0},\cdots,\hat{\mathbf{b}}_{\hat{N}}\}\in\mathbb{R}^{d_{b}}$ similar to $\mathbf{S}$, but with the dynamics reduced to a velocity vector. Based on the 3D bounding boxes, as well as a rasterized bird’s-eye view (BEV) map $\mathcal{M}\in\mathcal{R}^{h\times w\times 5}$ containing information on the static scene layout, the planner then computes a future trajectory $\boldsymbol{\tau}$ to be driven.
The detector incurs errors in the form of false-positives, false-negatives and inaccurate regression of bounding box parameters when compared to the ground-truth set derived directly from $\mathbf{S}$. These errors are individual to the specific model architecture and weights in question, and often depend on the configuration of the scene, e.g. due to correlations with the relative object pose, correlations among groups of objects or correlations between objects and map elements. To probe the planner’s sensitivity to this imperfect perception, we would like to find worst-case, but plausible errors for a given scene that drive the planned trajectory towards violation of safety constraints. To this end, we learn a conditional generative PEM as a proxy that imitates the detector given the ground-truth scene context $(\mathbf{S},\mathcal{M})$ as input. By repeatedly drawing samples $\tilde{\mathbf{B}}=\{\tilde{\mathbf{b}}_{0},\cdots,\tilde{\mathbf{b}}_{\tilde{N}}\}\in\mathbb{R}^{d_{b}}$ from this proxy and measuring the quality of the corresponding planned trajectory, we aim to gauge the influence of noise patterns and optimize for planning failures, such as a collision.
@@ III-A Conditional Generative PEM
We cast the imitation of perception errors given a ground-truth scene state as a set-to-set modeling problem, for which the transformer architecture is a natural choice due to its permutation-invariant modeling of relationships among set elements. We propose a conditional generative model building on a transformer encoder-decoder architecture [27] under the framework of CVAEs [28]. Specifically, we aim to construct a latent variable model capturing
@@
where $\mathbf{Z}=\{\mathbf{z}_{0},\cdots,\mathbf{z}_{\hat{N}}\}\in\mathbb{R}^{d_{z}}$ is a set of per-object latent variables of dimensionality $d_{z}$ explaining the stochasticity of the target detector’s perception errors by capturing different noise patterns. This allows sampling multiple plausible sets of detections $\tilde{\mathbf{B}}\sim p_{\theta}(\hat{\mathbf{B}}\mid\mathbf{S},\mathcal{M})$ by applying different latent error characteristics sampled from a learned conditional Gaussian prior modeling $\mathbf{Z}\sim p_{\theta}(\mathbf{Z}\mid\mathbf{S},\mathcal{M})$ to a given scene context $(\mathbf{S},\mathcal{M})$. This is done via a deterministic decoder $\hat{\mathbf{B}}=f_{\theta}(\mathbf{Z},\mathbf{S},\mathcal{M})$. As evaluating the integral in (1) is intractable, learning under the CVAE framework utilizes a learned approximation to the posterior $q_{\phi}(\mathbf{Z}\mid\mathbf{S},\mathcal{M},\hat{\mathbf{B}})$ to maximize the evidence lower bound (ELBO[29, 28]:
@@
Although both the prior and the posterior distribution are factorized over objects, the distribution over each latent variable $\mathbf{z}_{n}\in\mathbf{Z}$ is computed from the full set of ground-truth states $\mathbf{S}$ (and detector outputs $\hat{\mathbf{B}}$ for the posterior). Similarly, the decoder considers the joint sets of per-object latent variables $\mathbf{Z}$ and states $\mathbf{S}$ to output individual detections $\tilde{\mathbf{b}}_{n}\in\tilde{\mathbf{B}}$. This enables our model to generate realistic, scene-consistent noise patterns by allowing it to capture higher-order relationships. In the following, we briefly describe the main components in our architecture, as depicted in Fig. 2.
@@
Decoder: Given the set of initial hypotheses and map context, the decoder applies a sampled latent error pattern $\mathbf{Z}$ to produce a set of proxy bounding boxes $\tilde{\mathbf{B}}=f_{\theta}(\mathbf{Z},\mathbf{S},\mathcal{M})$. To this end, the information in both $\mathbf{Q}_{H}$ and $\mathbf{Z}$ is first fused by applying an MLP to the feature-wise concatenation of each pair of vectors $(\mathbf{q}_{n},\mathbf{z}_{n})$, forming $\tilde{\mathbf{Q}}_{H}\in\mathcal{R}^{d_{h}}$. Similarly to the prior encoder, $\tilde{\mathbf{Q}}_{H}$ is then refined through a series of $L_{d}$ transformer layers performing self-attention and cross-attending to $\mathbf{Q}_{\mathcal{M}}$. Finally, the refined latent queries are decoded to the bounding box regression parameters and independent per-class sigmoid classification scores. This mirrors the prevalent output configuration used in state-of-the-art 3D detectors [2, 30, 3].
Training: We optimize the model on a dataset of tuples $(\mathbf{S},\mathcal{M},\hat{\mathbf{B}})$ collected by running inference with a target detector on the corresponding sensor data to obtain noisy detections. Similar to prior work [31, 25], we use a modified CVAE objective $\mathcal{L}_{\text{cvae}}=\mathcal{L}_{\text{recon}}+\beta\mathcal{L}_{\text{JS}}$ consisting of a reconstruction error and a divergence regularizer:
@@ III-B Imitation Learning-based Planner
We consider a simple transformer-based planner that is prototypical of the planning modules proposed in recent literature [7, 35, 36, 37]. It operates on BEV projections of $\hat{\mathbf{B}}$ or $\tilde{\mathbf{B}}$, which are encoded via an MLP to form $\mathbf{Q}_{H}^{\pi}\in\mathbb{R}^{d_{\pi}}$, as well as a set of map features $\mathbf{Q}_{\mathcal{M}}^{\pi}\in\mathbb{R}^{d_{\pi}}$ obtained from a convolutional encoder. The planner then forms a latent plan by refining an initial constant learnable embedding $\mathbf{q}_{\text{ego}}^{\pi}$ through a series of $L_{\pi}$ transformer layers consisting of two cross-attention blocks attending to the encoded detection results $\mathbf{Q}_{H}^{\pi}$ and map features $\mathbf{Q}_{\mathcal{M}}^{\pi}$. To provide additional context, we add embeddings of the current speed of the ego vehicle $\vartheta^{\text{ego}}$, as well as a high-level navigation command $c^{\text{nav}}$ (i.e. go-straight, turn-left or turn-right) to the initial $\mathbf{q}_{\text{ego}}^{\pi}$. Finally, an MLP decodes a trajectory $\hat{\boldsymbol{\tau}}=\pi_{\omega}(\mathbf{B},\mathcal{M},\vartheta^{\text{ego}},c^{\text{nav}})\in\mathbb{R}^{T_{\text{plan}}\times 3}$ of future waypoints consisting of a BEV position and heading angle. The model is trained via standard imitation-learning on expert trajectories using an $l_{1}$-loss.